the tickets for a dance recital cost $5.00 for adults and $2.00 for children. if the total number of tickets sold was 295 and the total amount collected was $1,220, how many adult tickets were sold? [use an algebraic solution.]
Answer by stanbon(75887)

the tickets for a dance recital cost $5.00 for adults and $2.00 for children. if the total number of tickets sold was 295 and the total amount collected was $1,220, how many adult tickets were sold? [use an algebraic solution.]
------------------------
Let "a" be # of adult tickets; Let "c" be # of child tickets.
EQUATIONS:
Quantity Eq.: a + c = 295
Value Eq....:5a + 2c = 1220
-----------------
Multiply 1st eq. by 2 to get:
2a + 2c = 590
------------
Subtract that from the 2nd equation to solve for "a":
3a = 630
a = 210 (# of adult tickets sold)
----------------
Substitute into a+c=295 to solve for "c":
210 + c = 295
c = 85 (# of child tickets sold)
=================
Cheers,
Stan H.